• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Saha

    Log entry time 16:02:18 on June 2,2004

    Entry number 124418

    keyword=Shift Summary

    Shift Summary

    0800: Hall in Controlled access. Techs in hall to fix right dipole.
    1500: Beam back. Septum magnets and Left HRS Dipole ramping up.
    1515: Left Dipole ramping down slowly. Tech called in.
    1530: Everything OK. Left Septum still ramping up.
    1600: Everything set to go The next shift is ready to take data.